OpenStudy (anonymous):

Solve (2/3)x - 4 = 2/4

OpenStudy (across):

we have \[\frac{2}{3}x-4=\frac{2}{4}\]isolate x \[\frac{2}{3}x=\frac{2}{4}+4\]\[\frac{2}{3}x=\frac{9}{2}\]\[x=\frac{27}{4}\]

jimthompson5910 (jim_thompson5910):

\[\large \frac{2}{3}x-4=\frac{2}{4}\] \[\large \frac{2}{3}x-4=\frac{1}{2}\] \[\large 6(\frac{2}{3}x-4)=6(\frac{1}{2})\] \[\large 4x-24=3\] \[\large 4x=3+24\] \[\large 4x=27\] \[\large x=\frac{27}{4}\]

OpenStudy (anonymous):

ty